# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ Abstract.
Differential privacy (DP) has been integrated into graph neural networks (GNNs) to protect sensitive structural information, e.g., edges, nodes, and associated features across various applications. A common approach is to perturb the message-passing process, which forms the core of most GNN architectures. However, existing methods typically incur a privacy cost that grows linearly with the number of layers (Usenix Security’23), ultimately requiring excessive noise to maintain a reasonable privacy level. This limitation becomes particularly problematic when deep GNNs are necessary to capture complex and long-range interactions in graphs.
@@
In this paper, we theoretically establish that the privacy budget can converge with respect to the number of layers by applying privacy amplification techniques to the message-passing process, exploiting the contractive properties inherent to standard GNN operations. Motivated by this analysis, we propose a simple yet effective Contractive Graph Layer (CGL) that ensures the contractiveness required for theoretical guarantees while preserving model utility. Our framework, $\mathsf{CARIBOU}$, supports both training and inference, equipped with a contractive aggregation module, a privacy allocation module, and a privacy auditing module. Experimental evaluations demonstrate that $\mathsf{CARIBOU}$ significantly improves the privacy-utility trade-off and achieves superior performance in privacy auditing tasks.
Differential Privacy, Convergent Privacy Analysis, Graph Neural Networks, Perturbed Message Passing, Privacy Amplification.
$\ast$ co-corresponding authors.
ccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Privacy protectionsccs: Computing methodologies Machine learning algorithms
1. Introduction
@@
Graph neural networks (GNNs) (Kipf and Welling, [n. d.]; Gu et al., 2020; Velickovic et al., 2018), designed for operating over structural data, have achieved success in various domains, including social networks (Perozzi et al., 2014; Zhang et al., 2022b) and recommendation systems (Chen et al., 2024a; Wu et al., 2021b, 2023a). At their core, many GNN architectures are built upon the message-passing paradigm, where node representations are iteratively updated by aggregating information from their neighbors. However, graph structures often encode sensitive information about relationships and attributes. As a result, GNNs — like other machine learning models — are vulnerable to privacy attacks, including membership inference (Wu et al., 2022a; Ye et al., 2022; He et al., 2021b) and attribute inference (Zhang et al., 2022a; Sun et al., 2023). These vulnerabilities highlight the urgent need for robust privacy protection mechanisms in graph learning.
@@
Differential privacy (DP) (Mironov, 2017; Wasserman and Zhou, 2010; Dong et al., 2022; Sun et al., 2024) has emerged as a foundational framework for providing formal guarantees against data leakage over the past two decades, with widespread applications in machine learning (Tang et al., 2023; Yu et al., 2019), synthetic data generation (Sun et al., 2024; Li et al., 2023a; McKenna et al., 2021; Patwa et al., 2023), and beyond. In the context of GNNs, recent works (Sajadmanesh and Gatica-Perez, 2020; Wu et al., 2023b) have advanced privacy protection through edge-level DP (EDP) (Wu et al., 2022b) and node-level DP (NDP) (Iftikhar and Wang, 2021; Day et al., 2016) guarantees. The primary approach employs perturbed message passing, which injects calibrated Gaussian or Laplace noise into aggregation layers to protect the edge or node memberships in the training graph. While these approaches have provided formal privacy guarantees, they share a critical limitation: the privacy loss grows linearly with the number of layers $K$ or graph hops. In other words, each additional aggregation layer compounds the privacy cost, ultimately requiring large amounts of noise to maintain a reasonable level of privacy protection. This, in turn, severely degrades model utility. This is particularly problematic when deep GNNs are needed for capturing complex relationships (Feng et al., 2022) or when analyzing graphs with long-range interactions (Dwivedi et al., 2022; Chen et al., 2024b; Wu et al., 2021a; Tönshoff et al., 2024; Thürlemann and Riniker, 2023), such as molecular graphs (Zhang et al., 2022c), as in these cases the message-passing mechanism requires multiple hops to reach distant nodes.
@@
Interestingly, empirical studies (Wu et al., 2022a) have shown that membership inference attacks are not particularly more successful against deeper GNNs, suggesting that the linear dependency of privacy cost on network depth might be an overestimation. This observation aligns with the phenomenon known as “over-smoothing” (Chen et al., 2020a) in GNNs, where node representations become increasingly homogeneous as network depth increases and consequently making membership inference more challenging. This homogenization effect might actually provide inherent privacy benefits due to the contractive nature of GNN aggregation operations. This challenge motivates our central research question:
@@
Can we achieve differentially private graph learning with a convergent (bounded) privacy budget, thereby improving the privacy-utility trade-off for deep GNNs?
@@
Figure 1. Comparing Calibrated Noise for Perturbed Message Passing. Previous analysis requires $\sigma^{2}\propto O(K)$, and our analysis demands $\sigma^{2}\propto O\Bigl{(}(1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})/((1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}}))\Bigl{)}$, where $C_{\mathsf{L}}$ is Lipchitz constant. With sensitivity constrained to norm $1$, the signal-to-noise ratio is markedly low as $K$ increases, severely impacting utility.
In this work, we answer this question affirmatively. Unlike previous studies, we propose a new privacy analysis that leverages the contractiveness of message-passing operations to achieve a convergent privacy budget with respect to the number of layers $K$ instead of a linear growth as in previous works; see Figure 1 for a comparison. We further design a novel privacy-preserving GNN framework, $\mathsf{CARIBOU}$, that incorporates this analysis into a practical model.
1.1. Overview of Convergent Privacy Analysis
@@
For a perturbed message-passing GNN with $K$ layers, the standard approach analyzes the privacy loss at each step and then applies the DP composition theorem to aggregate the total privacy cost. This approach is common in existing privacy analyses of perturbed message-passing GNNs (Sajadmanesh and Gatica-Perez, 2020; Wu et al., 2023b; Chien et al., 2023), resulting in a privacy budget that scales linearly with $K$, specifically $\epsilon=O(K/\sigma^{2})+O(\sqrt{K/\sigma})$. Thus, as $K$ increases, the amount of injected noise $\sigma$ must also grow, leading to degraded utility — particularly when a small $\epsilon$ is desired; see Section 3 for a motivating empirical study.
@@
Inspired by recent advances in privacy amplification (Chourasia et al., 2021; Bok et al., 2024) through hidden states and contractive iterative processes, we observe that a similar amplification effect can be exploited in GNNs. Here, contractiveness refers to the property that the distance between two inputs is reduced after applying the operation, implying reduced distinguishability of outputs from a privacy perspective. The potential privacy amplification in GNNs arises from the following two observations: (1) GNNs typically do not expose intermediate node embeddings during training or inference, focusing only on the final node representations. (2) Standard message-passing operations — such as those used in Graph Convolutional Networks (GCNs) (Kipf and Welling, [n. d.]) (the dominant model in practice and in empirical studies (Wu et al., 2022a)) — are inherently contractive, a property that also underlies the over-smoothing phenomenon (Chen et al., 2020a). We theoretically validate this insight by showing that the privacy loss of a $K$-layer perturbed message-passing process with contractive layers satisfies a convergent privacy budget. Specifically, instead of growing linearly with $K$, we show that the privacy budget follows a convergent form: $\epsilon=O\left(\frac{(1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})}{(1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}})}\right),$ where $C_{\mathsf{L}}$ is the Lipschitz constant of the message-passing operator; see Theorem 6 for details. Our improved privacy analysis is achieved by recasting the multi-layer perturbed GNN as a Contractive Noisy Iteration (CNI) process (Bok et al., 2024) and applying the privacy convergence results established for CNIs.
@@
To leverage this analysis in practice, we design a simple yet effective Contractive Graph Layer (CGL) that enforces the contractiveness required for our theoretical guarantees while maintaining model expressivity. The CGL layer builds upon standard GCN-style aggregation, augmented with residual connections (Chen et al., 2020b) and mean aggregation normalization (Alon and Yahav, 2020), ensuring expressiveness across many layers without exposing sensitive edge information.
@@
1.2. $\mathsf{CARIBOU}$: Framework and Evaluation
@@
To evaluate $\mathsf{CARIBOU}$, we conduct extensive experiments over nine graph datasets, including commonly-used real-world datasets and chain-structured datasets for developing configurable interaction ranges. The experimental results demonstrate that $\mathsf{CARIBOU}$ improves non-trivial utility over standard graph and chain-structured datasets. Compared with several SOTA baselines, $\mathsf{CARIBOU}$ shows significant utility improvements, especially in high privacy regimes, and reasonable computational overhead. Ablation studies are provided to understand the relation between privacy-utility hops and various ranges of graph, and the choice of hyper-parameters of CGL. In addition to privacy verification, we perform auditing experiments based on two membership inference attacks (Wu et al., 2022a; Olatunji et al., 2021b) , showing $\mathsf{CARIBOU}$’s resistance to attacks.
Contribution. To summarize, our contribution includes:
  1. (1)
@@
    A novel privacy analysis for GNNs that leverages the contractiveness of message-passing operations to achieve convergent privacy costs, even for deep networks; (Section 5)
  2. (2)
@@
    The design of perturbed CGL and a practical differentially private GNN framework – $\mathsf{CARIBOU}$ with provable privacy guarantees and superior utility-privacy tradeoffs; (Section 4)
  3. (3)
@@
    Extensive experimental validation across multiple graph datasets with varying structural properties, demonstrating significant improvements over state-of-the-art private GNN approaches. (Section 6)
1.3. Related Works
DP for graphs.
Graph can be interpreted into the combination by edges and nodes. Corresponding to instance-level DP (Mironov, 2017; Abadi et al., 2016; Ye and Shokri, 2022), the “instance” of graph data can be an edge or a node, naturally called edge DP and node DP. Early stage works (Gehrke et al., 2011; Mir and Wright, 2009; Wang et al., 2013; Chen et al., 2014; Lu and Miklau, 2014) started by private statistics estimation or counting, or release of private graphs under edge DP. Node DP protects against revealing the presence or absence of an individual node along with all its adjacent edges. Prior works consider node-privacy (Kasiviswanathan et al., 2013; Iftikhar and Wang, 2021; Day et al., 2016; Xia et al., 2021; Imola et al., 2021) tailored to specific graph statistics constrained by bounded node degrees. Designing DP mechanisms for graphs requires thoughtful consideration of multi-hop aggregation, demanding careful design to balance privacy and utility.
Differentially private GNNs.
GNNs have emerged as a key approach for applications over graph-structured data, such as intrusion detection (Apruzzese et al., 2023; Sommer and Paxson, 2010), social recommendation (Xia et al., 2023), and drug discovery (Zhong and Mottin, 2023). Sharing finally established model can lead to privacy risks (Zhang et al., 2022a; Ye et al., 2022; He et al., 2021b), typically membership inference attack (MIA) (Carlini et al., 2022; Ye et al., 2022). MIA stems from “overfitting”, where models can memorize training memberships (Zhang et al., 2021). In GNNs, message-passing mechanisms can leak sensitive information about their edge- or node-level neighbors. To address these risks, existing research (Sajadmanesh and Gatica-Perez, 2020; Wu et al., 2023b; Chien et al., 2023; Daigavane et al., 2021; Olatunji et al., 2021a; Sajadmanesh et al., 2023) integrating DP with GNNs, advance the EDP or NDP by a perturbed message-passing mechanism that adds the calibrated Gaussian (or Laplace) noise to the message-passing layer. Albeit their success, leveraging the contractive hidden node embeddings in private GNNs for enhancing privacy analysis remains an underexplored avenue. More related works are detailed in Appendix F.
2. Preliminary
2.1. Message Passing Graph Neural Networks
@@
Graph Neural Networks (GNNs) are a class of neural networks that operate on graph-structured data. Most GNNs follow the message-passing paradigm (Feng et al., 2022), where nodes iteratively aggregate information from their neighbors to update their representations.
Message Passing Layers.
@@
(1) $\text{MP}_{G}({\bm{X}}_{u}^{(k)}):=\sigma\left(\psi\left({\bm{X}}^{(k)},\oplus_{v\in{\mathcal{N}}(u)}\phi({\bm{X}}_{u}^{(k)},{\bm{X}}_{v}^{(k)})\right)\right),$
where $\sigma$ is a non-linear activation function, ${\mathcal{N}}(u)$ is the set of neighbors of node $u$, $\phi$ is a function that computes the message from node $v$ to node $u$, $\oplus$ represents the aggregation function that processes all messages from the neighbors of node $u$, and $\psi$ is a function that updates the node feature vector of node $u$ with the aggregated messages. Examples of message passing GNNs, such as GCN (Kipf and Welling, 2017) and its variant, are in Appendix A.
Applications of Message Passing GNNs.
Message passing GNNs leverage multiple layers to iteratively refine node representations, which are then employed in tasks like node classification (Kolluri et al., 2022), link prediction (Rao et al., 2024), and graph classification (Li et al., 2024). Long-range graphs (Dwivedi et al., 2022; Chen et al., 2024b; Tönshoff et al., 2024) enhance deeper GNNs (Thürlemann and Riniker, 2023; Wu et al., 2021a) by capturing dependencies between distant nodes, crucial for tasks like molecular property prediction (Ma et al., 2022), protein interaction modeling (Zhong et al., 2022), and model complex node interactions (Feng et al., 2022).
2.2. Differential Privacy for GNNs
Definition 0 (Differential Privacy (Dwork et al., 2006)).

### v2

# Convergent Privacy Framework for Multi-layer GNNs through Contractive Message Passing

@@ Abstract
Differential privacy (DP) has been integrated into graph neural networks (GNNs) to protect sensitive structural information, e.g., edges, nodes, and associated features across various applications. A prominent approach is to perturb the message-passing process, which forms the core of most GNN architectures. However, existing methods typically incur a privacy cost that grows linearly with the number of layers (e.g., GAP published in Usenix Security’23), ultimately requiring excessive noise to maintain a reasonable privacy level. This limitation becomes particularly problematic when multi-layer GNNs, which have shown better performance than one-layer GNN, are used to process graph data with sensitive information.
@@
In this paper, we theoretically establish that the privacy budget converges with respect to the number of layers by applying privacy amplification techniques to the message-passing process, exploiting the contractive properties inherent to standard GNN operations. Motivated by this analysis, we propose a simple yet effective Contractive Graph Layer (CGL) that ensures the contractiveness required for theoretical guarantees while preserving model utility. Our framework, $\mathsf{CARIBOU}$, supports both training and inference, equipped with a contractive aggregation module, a privacy allocation module, and a privacy auditing module. Experimental evaluations demonstrate that $\mathsf{CARIBOU}$ significantly improves the privacy-utility trade-off and achieves superior performance in privacy auditing tasks.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 24-28 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.240255 www.ndss-symposium.org §§footnotetext: Co-corresponding authors.
I Introduction
@@
Graph neural networks (GNNs) [iclr/KipfW17, nips/GuC0SG20, iclr/VelickovicCCRLB18], designed for operating over structural data, have achieved success in various domains, including social networks [kdd/PerozziAS14, www/ZhangGPH22] and recommendation systems [www/ChenBSXZHHWH24, sigir/WuWF0CLX21, csur/WuSZXC23]. At their core, many GNN architectures are built upon the message-passing paradigm, where node representations are iteratively updated by aggregating information from their neighbors. However, graph structures often encode sensitive information about relationships and attributes. As a result, GNNs are vulnerable to privacy attacks, including membership inference [sp/Wulinkteller22, ccs/0001MMBS22, corr/abs-2102-05429] and attribute inference [uss/000100S022, tkde/SunDYZWYHL23]. These vulnerabilities highlight the urgent need for robust privacy protection mechanisms in graph learning.
@@
Differential privacy (DP) [csf/mironov2017renyi, wasserman2010statistical, journal/dong2022gaussian] has emerged as a foundational framework to provide formal guarantees against data leakage over the past two decades, with widespread applications in machine learning [nips/TangPSM23, sp/Yu0PGT19, nips/YangZZ0PL023], synthetic data generation [imc/SunCGWL24, iclr/LiWC23, jpc/McKennaMS21, pvldb/PatwaSGMR23], and beyond. In the context of GNNs, recent works [arxiv/sajadmanesh2020differential, arxiv/Wuprivacy23] have advanced privacy protection through edge-level DP (EDP) [sp/0011L0022] and node-level DP (NDP) [pakdd/IftikharW21, sigmod/DayLL16] guarantees. The primary approach employs perturbed message passing, which injects calibrated Gaussian or Laplace noise into aggregation layers to protect the edge or node memberships in the training graph. While these approaches have provided formal privacy guarantees, they share a critical limitation: the privacy loss grows linearly with the number of layers $K$ or graph hops. In other words, each additional aggregation layer compounds the privacy cost, ultimately requiring large amounts of noise to maintain a reasonable level of privacy protection. This, in turn, severely degrades model utility.
Recent advances have shown that multi-layer GNNs, especially deeper GCNs, are essential in capturing complex relationships [feng2022powerful] and analyzing graphs with long-range interactions [nips/DwivediRGPWLB22, corr/abs-2406-03386, nips/WuJWMGS21, tmlr/TonshoffRRG24, iclr/ThurlemannR23]. In fact, informative long-range interactions exist in a lot of real-world datasets. For instance, in large biological networks, long-range dependencies influence protein functions, requiring more than 10 hops of message passing [icml/Li0GK21]. In social networks, privacy-sensitive relationships propagate through multi-hop neighborhoods [nguyen2017probability, liu2014assessment]. As reported in [icml/Li0GK21], increasing the network depth leads to a substantial improvement in accuracy from $72.5\%$ to $88.2\%$. However, the aforementioned linear dependence on $K$ is particularly challenging for multi-layer GNNs as larger $K$ leads to larger privacy parameter $\epsilon$, a.k.a weak privacy guarantee.
@@
Interestingly, empirical studies [sp/Wulinkteller22] have shown that membership inference attacks are not particularly more successful against multi-layer GNNs, suggesting that the linear dependency of privacy cost on network depth might be an overestimation. This observation aligns with the phenomenon known as “over-smoothing” [aaai/ChenLLLZS20] in GNNs, where node representations become increasingly homogeneous as network depth increases and consequently making membership inference more challenging. This homogenization effect might actually provide inherent privacy benefits due to the contractive nature of GNN aggregation operations. This observation motivates our central research question:
@@
Can we achieve differentially private graph learning with a convergent (bounded) privacy budget, thereby improving the privacy-utility trade-off for deeper GNNs?
@@
Figure 1: Comparing Calibrated Noise for Perturbed Message Passing. Previous analysis requires $\sigma^{2}\propto O(K)$, and our analysis demands $\sigma^{2}\propto O\Bigl((1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})/((1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}}))\Bigl)$, where $C_{\mathsf{L}}$ is Lipchitz constant. With sensitivity constrained to norm 11, the signal-to-noise ratio is markedly low as $K$ increases, severely impacting utility.
In this work, we answer this question affirmatively. Prior perturbed message-passing mechanisms [sp/0011L0022, pakdd/IftikharW21, sigmod/DayLL16] assume that privacy loss grows linearly with depth, yet empirical results show that deeper GNNs can be less vulnerable to membership inference. We attribute this phenomenon to the contractive nature of common aggregation operators. In theory, we analyze this contractive property underlying over-smoothing, which leads to bounded sensitivity, so the privacy budget converges with the number of layers $K$ (see Figure 1) instead of growing linearly. This motivates $\mathsf{CARIBOU}$, a privacy-preserving GNN framework that enforces contractiveness to mirror real-world GNN behaviors while achieving convergent privacy budget through privacy amplification.
I-A Overview of Convergent Privacy Analysis
@@
For a perturbed message-passing GNN with $K$ layers, the standard approach analyzes the privacy loss at each step and then applies the DP composition theorem to aggregate the total privacy cost. This approach is common in existing privacy analyses of perturbed message-passing GNNs [arxiv/sajadmanesh2020differential, arxiv/Wuprivacy23, nips/chien2023differentially], resulting in a privacy budget that scales linearly with $K$, specifically $\epsilon=O(K/\sigma^{2})+O(\sqrt{K/\sigma})$. Thus, as $K$ increases, the amount of injected noise $\sigma$ must also grow, leading to degraded utility particularly when a small $\epsilon$ is desired; see Section III for a motivating empirical study.
@@
Inspired by recent advances in privacy amplification [nips/ChourasiaYS21, icml/BokShifted24] through hidden states and contractive iterative processes, we observe that a similar amplification effect can be exploited in GNNs. Here, contractiveness refers to the property that the distance between two inputs is reduced after applying the operation, implying reduced distinguishability of outputs from a privacy perspective. The potential privacy amplification in GNNs arises from the following two observations: (1) GNNs typically do not expose intermediate node embeddings during training or inference, focusing only on the final node representations. (2) Standard message-passing operations, such as those used in Graph Convolutional Networks (GCNs) [iclr/KipfW17] (the dominant model in practice and in empirical studies [sp/Wulinkteller22]), are inherently contractive, a property that also underlies the over-smoothing phenomenon [aaai/ChenLLLZS20]. We theoretically validate this insight by showing that the privacy loss of a $K$-layer perturbed message-passing process with contractive layers satisfies a convergent privacy budget. Specifically, instead of growing linearly with $K$, we show that the privacy budget follows a convergent form: $\epsilon=O\left(\frac{(1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})}{(1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}})}\right),$ where $C_{\mathsf{L}}$ is the Lipschitz constant of the message-passing operator (see Theorem 3 for details). Our improved privacy analysis is achieved by recasting the multi-layer perturbed GNN as a Contractive Noisy Iteration (CNI) process [icml/BokShifted24] and applying the privacy convergence results established for CNIs.
@@
To leverage this analysis in practice, we design a simple yet effective Contractive Graph Layer (CGL) that enforces the contractiveness required for our theoretical guarantees while maintaining model expressivity. The CGL layer builds upon standard GCN-style aggregation, augmented with residual connections [icml/ChenWHDL20] and mean aggregation normalization [alon2020bottleneck], ensuring expressiveness across many layers without exposing sensitive edge information.
@@
I-B $\mathsf{CARIBOU}$: Framework and Evaluation
@@
To evaluate $\mathsf{CARIBOU}$, we conduct extensive experiments over nine graph datasets, including commonly-used real-world datasets and synthetic chain-structured datasets for developing configurable interaction ranges. The experimental results demonstrate that $\mathsf{CARIBOU}$ improves non-trivial utility over standard graph and chain-structured datasets. Compared with several SOTA baselines, $\mathsf{CARIBOU}$’s EDP and NDP show significant utility improvements, especially in high privacy regimes, and reasonable computational overhead. Table I presents a comprehensive comparison, which is explained in Section VIII. Ablation studies are provided to understand the relation between privacy-utility hops and various ranges of graph, and the choice of hyper-parameters of CGL. In addition to privacy verification, we perform auditing experiments based on two membership inference attacks [sp/Wulinkteller22, tpsisa/OlatunjiNK21], demonstrating $\mathsf{CARIBOU}$’s robustness.
Contribution. In terms of our new insights (Section III), our contribution includes:
  1. 1.
@@
    A novel privacy analysis for GNNs that leverages the contractiveness of message-passing operations to achieve convergent privacy costs, even for deep networks; (Section V)
  2. 2.
@@
    The design of perturbed CGL and a practical differentially private GNN framework – $\mathsf{CARIBOU}$ with provable privacy guarantees and superior utility-privacy tradeoffs; (Section IV)
  3. 3.
@@
    Extensive experimental validation across multiple graph datasets with varying structural properties, demonstrating significant improvements over state-of-the-art private GNN approaches. (Section VI)
TABLE I: Comparison between Private GNNs. EDP and NDP summarizes the results of private GNNs in Table III.
Framework Mechanism Complexity per Layer Calibrated Noise ($\sigma$) EDP Utility NDP Utility PertGraph [ccs/KolluriBHS22, sp/0011L0022] Graph perturbation $O(|{\bm{V}}|^{2})$ $\propto 1$ DPDGC [nips/chien2023differentially] Decoupled graph with perturbation $O(|{\bm{E}}|)$ $\propto\sqrt{K}$ GAP [uss/sajadmanesh2023gap] Perturbed message passing $O(|{\bm{E}}|)$ $\propto\sqrt{K}$ $\mathsf{CARIBOU}$ Perturbed message passing $O(|{\bm{E}}|)$ $\propto\sqrt{\min(K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}})}$
II Preliminary
II-A Message Passing Graph Neural Networks
@@
Graph Neural Networks (GNNs) are a class of neural networks that operate on graph-structured data. Most GNNs follow the message-passing paradigm [feng2022powerful], where nodes iteratively aggregate information from their neighbors to update their representations.
II-A1 Message Passing Layers
@@
$\text{MP}_{G}({\bm{X}}_{u}^{(k)}):=\sigma\left(\psi\left({\bm{X}}^{(k)},\oplus_{v\in{\mathcal{N}}(u)}\phi({\bm{X}}_{u}^{(k)},{\bm{X}}_{v}^{(k)})\right)\right),$ (1)
where $\sigma$ is a non-linear activation function, ${\mathcal{N}}(u)$ is the set of neighbors of node $u$, $\phi$ is a function that computes the message from node $v$ to node $u$, $\oplus$ represents the aggregation function that processes all messages from the neighbors of node $u$, and $\psi$ is a function that updates the node feature vector of node $u$ with the aggregated messages. Examples of message passing GNNs, such as GCN [iclr/kipf2017semi] and its variant, are in Appendix -A.
II-A2 Applications of Message Passing GNNs
Message passing GNNs leverage GNN layers to iteratively refine node representations, which are then employed in tasks like node classification [ccs/KolluriBHS22], link prediction [www/RaoKR24], and graph classification [icdm/LiW024]. Multi-layer GNNs like deep GNNs [iclr/ThurlemannR23, nips/WuJWMGS21] are especially suitable to process long-range graphs [nips/DwivediRGPWLB22, corr/abs-2406-03386, tmlr/TonshoffRRG24] by capturing dependencies between distant nodes, which is crucial for tasks like molecular property prediction [bioinformatics/MaBRHXXYH22], protein interaction modeling [bmcbi/ZhongHXLQY22], and complex node interaction modeling [feng2022powerful].
II-B Differential Privacy for GNNs
Definition 1 (Differential Privacy [tcc/DworkMNS06]).